A new father plans on saving for his daughter’s college education.  He will donate $1,000 on her first birthday.  After that, he will increase his donation by 4.50% each year and will make his last contribution on her 18th birthday.  If he can earn 7.00% each year in his investment account, how much will his daughter’s college fund be worth on her 18th birthday?
Answer #1

Donation on first Birthday = $1000

His donation will increase by 4.50% each year till the last contribution on 18th birthday.

Calculating the Future Value of these Donation:-

FV of Growing annuity = C*\frac{[(1+r)^{n}-(1+g)^{n}]}{r-g}

Where, C= Cash Flow = $1000

r = Periodic Interest rate = 7%

g = Growth rate of each cash flow = 4.50%

n= no of periods =18

FV = 1000*\frac{[(1+0.07)^{18}-(1+0.045)^{7}]}{(0.07-0.045)}

FV = 1000*\frac{[3.37993227573-2.20847876636]}{(0.025)}​​​​​​

Future Value = $46,858.14

So, Balance in daughter's college fund at her 18th age is $46,858.14

If you need any clarification, you can ask in comments.     

If you like my answer, then please up-vote as it will be motivating